40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 16 18:08:57 2016, Johannes, Update, PSL, Before the AOM installation IMG_20160615_154556392e.jpg
    Reply  Fri Jun 17 22:36:11 2016, Johannes, Update, PSL, AOM installation aom_new_mount.jpgaom_beam_split.jpgaom_dump_IR.jpg
       Reply  Mon Jun 20 08:26:56 2016, Steve, Update, PSL, AOM pictures AOMinplace.jpginplaceAOM.jpgAOMin.png
Message ID: 12192     Entry time: Thu Jun 16 18:08:57 2016     Reply to this: 12196
Author: Johannes 
Type: Update 
Category: PSL 
Subject: Before the AOM installation 

There was only one razor blade beam dump labeled for atmospheric use left, but that's all we need. Steve is working on restocking. I placed the modified AOM mount on the PSL table near its intended location (near the AOM where it doesn't block any beams).

Things to keep in mind:

  • The laser power needs to be turned down for the installation of the AOM. Current laser settings are: Crystal Temperature: 29.41 C, Diode Current: 2.1 A.
  • The AOM driver must not be left unterminated when turned on (which it currently is and will be).
  • The HEPA filters are currently running at ~50%. While the PSL enclosure is open for the work we'll set them to 100% and lower them after a job well done.

The setup:

The AOM has a deflection angle of about 20 mrad, which requires about 10cm of path for a separation of 2mm of the two beams. I need to survey closer and confirm, but I hope I can fit the beam dump in before the PMC (this of course also depends on the spot size). Alternatively, the PMC hopefully isn't resonant for anything remotely relevant at 80MHz offset, in which case we can also place the beam dump in its reflection path.

So this is the plan:

  • Determine coupling efficiency into PMC for reference
  • Turn down laser power
  • No signal on AOM driver modulation input
  • Mount AOM, place in beam path, and align
  • Correct alignment into PMC?
  • Secondary beam detectable? Adjust modulation input and laser power until detectable.
  • Find a place for beam dump
  • Confirm that primary beam is not clipping with PMC
  • Turn up laser power
  • Determine coupling efficiency with restored power to compare

Any thoughts? Based on the AOMs resting place I assumed that it is supposed to be installed before the PMC, but I'm actually not entirely sure where it was sitting before.

ELOG V3.1.3-